OpenStudy (anonymous):

Determine the equation of the line that passes through the points (2, -5) and (7, 0). Use the form y = mx + b.

OpenStudy (anonymous):

is this a question?

OpenStudy (anonymous):

y = 1x - 7

OpenStudy (anonymous):

m = slope = rise (delta y) / run (delta x) m = (0 - -5) / (7 - 2) = 5 / 5 = 1 b = y intercept = value for y when x = 0; can be computed by using one of the know points (I will use (7,0) 0 = 1 * 7 + b 0 = 7 + b -7 = b By substitution, y = 1x -7 or y = x -7

OpenStudy (anonymous):

ahh, algebra!